# FuSeFL: Fully Secure and Scalable Cross-Silo Federated Learning

### v1

@@ 5.4 Communication Complexity
Server-to-Client / Client-to-Server (S2C/C2S). In AriaNN-FL, the server actively participates in secure training by running MPC protocols jointly with each client across every local epoch. This introduces a severe server-side bottleneck as the number of clients grows. In contrast, FuSeFL decouples training from the server and restricts its role to model aggregation. Clients only communicate their secret-shared updates to two non-colluding servers once per global round. The overall S2C/C2S complexity is reduced to $\mathcal{O}((n/2)\cdot x)$, assuming a model size of $x$ and $n$ total clients. This significantly improves scalability and enables parallel execution. Client-to-Client (C2C). AriaNN-FL performs client-to-client communication during aggregation, where each client sends its model share to a designated aggregator for reconstruction—incurring $\mathcal{O}(n\cdot x)$ communication and requiring one trusted aggregator per round. In contrast, FuSeFL confines secure training to fixed client pairs using MPC. Each pair exchanges secret shares of data and model parameters locally, with communication parallelized across groups. This keeps per-group C2C overhead constant and independent of the total number of clients.
By restructuring the communication pattern to exploit pairwise training and two-server aggregation, FuSeFL substantially reduces both the volume and centralization of communication (evaluated in §7.3). This architectural change eliminates per-epoch server interaction, enables system-wide parallelism, and improves bandwidth efficiency without weakening security guarantees.
5.5 The Memory Complexity
@@
5.6 Security Guarantees
FuSeFL provides strong security guarantees for both data and model confidentiality in a distributed, partially untrusted environment. These guarantees are realized through a combination of secret sharing, secure multi-party computation (MPC), and adversary-aware system design, as defined in our threat model (Section 3). Below, we detail the key properties and the cryptographic principles underpinning them.
@@ 5.7 Flexibility and Generalization
To maintain end-to-end confidentiality, the aggregation infrastructure must scale accordingly. Specifically, FuSeFL requires one aggregation server per client in the group to receive and process the respective secret shares. As a result, for a group of $n$ clients using an $n$-party MPC protocol, FuSeFL can tolerate up to $n{-}1$ colluding clients or servers without compromising data or model confidentiality.
@@ 6 Evaluation Methodology
Experiments are conducted on a machine equipped with an Intel(R) Xeon(R) W-2123 CPU (3.60 GHz, 4 cores). Our FuSeFL implementation is built on top of AriaNN’s implementation [22], which performs secure training using two workers while keeping the data and model private. In FuSeFL, the AriaNN training protocol is applied within each client group to collaboratively train the local model.
@@ 7.1 Accuracy Evaluation
Importantly, FuSeFL-Serial achieves comparable accuracy 1.43×1.43$\times$1.43 × faster than AriaNN-FL and 6.85×6.85$\times$6.85 × faster than AriaNN. For Network2 and LeNet, FuSeFL-Serial matches AriaNN’s accuracy—98.40% and 99.20%, respectively—demonstrating that FuSeFL delivers the same learning quality as traditional secure training, while offering substantial speedups due to its decentralized design. As observed in Network1, AriaNN-FL and FuSeFL-Parallel yield slightly lower accuracy for Network2 and LeNet, reaching 97.80% and 99.00%, respectively.
@@
7.2 Server-Side Scalability Analysis
@@
As illustrated in Fig. 9, AriaNN-FL’s training time per epoch grows significantly with the number of clients, due to its centralized training coordination. For example, with 128 clients, FuSeFL-Serial achieves up to a 5.65×5.65$\times$5.65 × speedup over AriaNN-FL, while FuSeFL-Parallel achieves up to 7.83×7.83$\times$7.83 ×. The bottleneck in AriaNN-FL arises from limited parallelism: if a server has 16 compute threads, it must process clients in sequential batches, e.g., training 64 clients in four rounds of 16. This serialization leads to latency that scales linearly with client count, severely limiting scalability.
@@
7.3 Communication Latency at Scale
@@ 7.5 Comparison with WW-FL[17]
TABLE VI: Training parameters of WW-FL and FuSeFL
Parameter WW-FL FuSeFL
@@
7.6 Robustness Under Threat Model
This section analyzes the robustness of FuSeFL under the adversarial threat model described in Section 3. The primary security goals are to protect the confidentiality of user data, model parameters, and intermediate updates in the presence of semi-honest adversaries. We consider two prominent threat vectors: unauthorized observation and model theft.
@@ 7.6 Robustness Under Threat Model
Protection Against Model Theft. Model confidentiality is a critical requirement in federated settings, particularly when the model itself represents proprietary intellectual property. In FuSeFL, the global model is secret-shared between two non-colluding servers and then transmitted to clients as encrypted shares. Each client receives only one share and cannot reconstruct the full model independently. Even if an adversary intercepts a share during transmission, it reveals no meaningful information.
@@ 8 Related Work
Other complementary works attempt to establish model ownership through watermarking [15, 16]. Backdoor watermarking uses a trigger dataset, while parameter watermarking embeds ownership directly in model weights. While effective for post-training verification, these methods do not provide confidentiality during training and are orthogonal to secure computation.

### v2

@@ V-D Communication Complexity
Server-to-Client / Client-to-Server (S2C/C2S). In AriaNN-FL, the server actively participates in secure training by running MPC protocols jointly with each client across every local epoch. This introduces a severe server-side bottleneck as the number of clients grows. In contrast, FuSeFL decouples training from the server and restricts its role to model aggregation. Clients only communicate their secret-shared updates to two non-colluding servers once per global round. The overall S2C/C2S complexity is reduced to $\mathcal{O}((n/2)\cdot x)$, assuming a model size of $x$ and $n$ total clients. This significantly improves scalability and enables parallel execution. Client-to-Client (C2C). AriaNN-FL performs client-to-client communication during aggregation, where each client sends its model share to a designated aggregator for reconstruction, which incurs $\mathcal{O}(n\cdot x)$ communication and requires one trusted aggregator per round. In contrast, FuSeFL confines secure training to fixed client pairs using MPC. Each pair exchanges secret shares of data and model parameters locally, with communication parallelized across groups. This keeps per-group C2C overhead constant and independent of the total number of clients.
By restructuring the communication pattern to exploit pairwise training and two-server aggregation, FuSeFL substantially reduces both the volume and centralization of communication (evaluated in §VII-C). This architectural change eliminates per-epoch server interaction, enables system-wide parallelism, and improves bandwidth efficiency without weakening security guarantees.
V-E The Memory Complexity
@@
V-F Security Guarantees
FuSeFL provides strong security guarantees for both data and model confidentiality in a distributed, partially untrusted environment. These guarantees are realized through a combination of secret sharing, secure multi-party computation (MPC), and adversary-aware system design, as defined in our threat model (Section III). Below, we detail the key properties and the cryptographic principles underpinning them.
@@ V-G Flexibility and Generalization
To maintain end-to-end confidentiality, the aggregation infrastructure must scale accordingly. Specifically, FuSeFL requires one aggregation server per client in the group to receive and process the respective secret shares. As a result, for a group of $n$ clients using an $n$-party MPC protocol, FuSeFL can tolerate up to n1n{-}1 colluding clients or servers without compromising data or model confidentiality.
@@ VI Evaluation Methodology
Experiments are conducted on a machine equipped with an Intel(R) Xeon(R) W-2123 CPU (3.60 GHz, 4 cores). Our FuSeFL implementation is built on top of AriaNN’s implementation [25], which performs secure training using two workers while keeping the data and model private. In our setup, both workers and the coordinating server run as separate processes on the same physical machine, communicating over a local network (LAN). In FuSeFL, the AriaNN training protocol is applied within each client group to collaboratively train the local model.
@@ VII-A Accuracy Evaluation
Importantly, FuSeFL-Serial achieves comparable accuracy $1.43\times$ faster than AriaNN-FL and $6.85\times$ faster than AriaNN. For Network2 and LeNet, FuSeFL-Serial matches AriaNN’s accuracy (98.40% and 99.20%, respectively), demonstrating that FuSeFL delivers the same learning quality as traditional secure training, while offering substantial speedups due to its decentralized design. As observed in Network1, AriaNN-FL and FuSeFL-Parallel yield slightly lower accuracy for Network2 and LeNet, reaching 97.80% and 99.00%, respectively.
@@
VII-B Server-Side Scalability Analysis
@@
As illustrated in Fig. 9, AriaNN-FL’s training time per epoch grows significantly with the number of clients, due to its centralized training coordination. For example, with 128 clients, FuSeFL-Serial achieves up to a $5.65\times$ speedup over AriaNN-FL, while FuSeFL-Parallel achieves up to $7.83\times$. The bottleneck in AriaNN-FL arises from limited parallelism: if a server has 16 compute threads, it must process clients in sequential batches, e.g., training 64 clients in four rounds of 16. This serialization leads to latency that scales linearly with client count, severely limiting scalability.
@@
VII-C Communication Latency at Scale
@@ VII-E Comparison with WW-FL[20]
TABLE VI: Training parameters of $W$-FL and FuSeFL
Parameter $W$-FL FuSeFL
@@
VII-F Robustness Under Threat Model
This section analyzes the robustness of FuSeFL under the adversarial threat model described in Section III. The primary security goals are to protect the confidentiality of user data, model parameters, and intermediate updates in the presence of semi-honest adversaries. We consider two prominent threat vectors: unauthorized observation and model theft.
@@ VII-F Robustness Under Threat Model
Protection Against Model Theft. Model confidentiality is a critical requirement in federated settings, particularly when the model represents proprietary intellectual property. In FuSeFL, the global model is secret-shared between two non-colluding servers and then transmitted to clients as encrypted shares. Each client receives only one share and cannot reconstruct the model independently. Even if an adversary intercepts a share during transmission, it reveals no meaningful information.
@@ VIII Related Work
Other works attempt to establish model ownership through watermarking [18, 19]. Backdoor watermarking uses a trigger dataset, while parameter watermarking embeds ownership directly in model weights. While effective for post-training verification, these methods do not provide confidentiality during training and are orthogonal to secure computation.